Can someone please help me, and let show me each step?
x-5[x-4x(x-3)]=24

Answer by stanbon(75887)

x-5[x-4x(x-3)]=24
----------------
Work form inside the brackets to outside the brackets:
---
x-5[x-4x^2+12x] = 24
----
x-5[-4x^2+13x] = 24
----
x +20x^2-65x = 24
---
20x^2-64x-24 = 0
----
5x^2-16x-6 = 0
---
x = [16 +- sqrt(16^2-4*5*-6)]/10
----
x = [16 +- sqrt(376)]/10
----
x = [16 +- 19.4]/10
============================
Cheers,
Stan H.
============